# Index Terms: Weakly supervised video anomaly detection, event prompt, multi-prompt learning, text.I Introduction

### v1

@@ Abstract
Most models for weakly supervised video anomaly detection (WS-VAD) rely on multiple instance learning, aiming to distinguish normal and abnormal snippets without specifying the type of anomaly. The ambiguous nature of anomaly definitions across contexts introduces bias in detecting abnormal and normal snippets within the abnormal bag. Taking the first step to show the model why it is anomalous, a novel framework is proposed to guide the learning of suspected anomalies from event prompts. Given a textual prompt dictionary of potential anomaly events and the captions generated from anomaly videos, the semantic anomaly similarity between them could be calculated to identify the suspected anomalous events for each video snippet. It enables a new multi-prompt learning process to constrain the visual-semantic features across all videos, as well as provides a new way to label pseudo anomalies for self-training. To demonstrate effectiveness, comprehensive experiments and detailed ablation studies are conducted on four datasets, namely XD-Violence, UCF-Crime, TAD, and ShanghaiTech. Our proposed model outperforms most state-of-the-art methods in terms of AP or AUC (82.6%, 87.7%, 93.1%, and 97.4%). Furthermore, it shows promising performance in open-set and cross-dataset cases.
@@
Weakly supervised video anomaly detection, event prompt, multi-prompt learning, text.I Introduction
Video anomaly detection (VAD) is crucial in video surveillance, given the extensive use of surveillance cameras. The task of VAD is to determine whether each frame in a video is normal or abnormal, which poses a significant challenge as it is not feasible to train a model with complete supervision. Consequently, weakly supervised learning methods (WS-VAD) [20, 62] that solely rely on video-level annotations have gained importance and popularity in recent years.
The general paradigm of these methods involves utilizing convolutional networks such as 3D ConvNet (C3D) [91], inflated 3D ConvNet (I3D) [93], or vision transformer [19] to extract visual features and aggregate spatio-temporal information between consecutive frames. Subsequently, an anomaly-detection network is trained using multiple instance learning (MIL) [96]. This approach simultaneously maximizes and minimizes the top- highest scores from individual anomaly and normal videos, respectively. However, most methods [96, 13] only considered the visual-related modality, and very few works [104] tried to introduce the semantic description of the videos. As a result, the MIL based approaches often suffer from a high false alarm rate (FAR) and low accuracy in detecting ambiguous abnormal events.
@@
The main contributions of this work are three-fold,
[leftmargin=2em]
To our best knowledge, this is the first work to introduce textual abnormal event prompts into video anomaly detection. The score predictor implicitly learns the definition of anomalies from the semantic features of anomaly prompts. It leads to incredible performance on open-set and cross-database problems.
A new multi-prompt learning strategy is proposed to provide an overall understanding of normal and abnormal patterns across different videos, while MIL is limited in individual videos.
@@ II-A Weakly Supervised Video Anomaly Detection
The weakly supervised video anomaly detection landscape primarily relies on multiple instance learning (MIL) due to limited annotated labels [96, 39]. However, conventional MIL faces challenges in providing sufficient supervision for various anomalies, leading to misclassifications and a high false alarm rate. To address these issues, Yu et al. propose cross-epoch learning (XEL) [20], which stores hard instances from previous epochs to optimize the anomaly predictor in the latest epoch. Additionally, UR-DMU [103] extends the anomaly memory unit into learnable dual memory units to alleviate the high false alarm rate issue. Another approach called robust temporal feature magnitude learning (RTFM) [13] trains a feature magnitude learning function to effectively recognize positive instances. All these methods are based on single or multiple visual modalities, including RGB and optical flow.
@@ II-B Prompt Tuning for Visual Tasks
In the realm of pretrained foundation multimodality models, a cost-effective prompt-tuning approach is gaining traction for adapting models to downstream tasks in natural language processing (NLP) [2, 3] and computer vision [4] domains.
The concept of prompt tuning originated in computer vision to tackle zero-shot or few-shot image tasks by incorporating semantic guidance. Multimodal models such as CLIP [109] leverage textual prompts for image classification, demonstrating state-of-the-art performance. In the video domain, Sato [6] explores prompt tuning for zero-shot anomaly action recognition, using skeleton features and text embeddings in a shared space to refine decision boundaries. Wang et al. introduce Prompt Learning for Action Recognition (PLAR) [7], which incorporates optical flow and learnable prompts to acquire both input-invariant knowledge from a prompt expert dictionary and input-specific knowledge based on the data.
A previous effort, PFMF [1], applies prompt-based learning to semi-supervised video anomaly detection. PFMF generates anomaly prompts by concatenating anomaly vectors from virtual datasets and scene vectors from real datasets, guiding the feature mapping network. However, the ”prompt” in PFMF defines anomalies at the visual level, introducing ambiguity. In our work, we propose textual anomaly prompts based on prior knowledge to mine fine-grained anomalies for achieving high performance.
III Proposed Method
@@ III-A Feature Synthesis
To clarify, we acquire the visual features and from video snippets in the abnormal and normal bags, utilizing the visual encoder of a CLIP model [109]. Given that many VAD videos, primarily from surveillance, often lack associated text descriptions, we leverage a pre-trained visual-to-text encoder from SwinBERT [111] to generate descriptions for each video snippet. These textual descriptions then undergo processing by the semantic feature extractor (SimCSE [110]), producing corresponding semantic features and for abnormal and normal video snippets.
@@
where symbolizes a feature alignment and fusion operation. Subsequently, the anomaly scores and can be calculated by applying and to a score predictor. Typically, this predictor takes the form of a multi-layer perceptron (MLP) [114], expressed as:
@@ III-B Multi-Prompt Learning
From Equation 5, it can be seen that the top- strategy focuses only on a few snippets with the highest scores within an individual video. Additionally, the top anomaly scores in an abnormal video may not be from an abnormal snippet. Therefore, a new textual prompt dictionary consisting of anomaly prompts is designed to link abnormal video snippets from different videos. For example, it could be ”A man is shooting a gun” or “Something is on fire”. As depicted in Figure 2, these prompts undergo the same semantic extraction process (SimCSE [110]) as the earlier video captions, generating their respective semantic features . Subsequently, we calculate the similarity between each prompt in the dictionary and every snippet in to construct an anomaly matrix as,
In essence, each element in provides insights into the probable type of anomaly associated with each snippet or indicates where a predetermined abnormal event might occur. Figure 3 offers a visual representation of , where frames containing abnormal events exhibit more pronounced colors. Notably, there is a discernible alignment between the frames and prompts.
@@ III-C Pseudo Anomaly Labeling
where and are the -th element of p and c, mean and std are the mean and standard deviation considering of the anomaly vector c. Then, the anomaly score predictor can be trained in a fully supervised manner, through a pseudo anomaly loss ,
@@
where and are hyper-parameters utilized in our model. Importantly, it’s worth noting that the MPL and PAL modules are trained collaboratively. During the inference stage, testing samples will only traverse the feature extractors and the predictor to acquire abnormal scores, with the MPL and PAL modules incurring no additional computational cost.
@@ IV Experiments
| Type | Source | Method | XD | UCF |
| AP(%) | AUC(%) | |||
| Semi | CVPR 16’ | Conv-AE [84] | 30.7 | - |
| CVPR 22’ | GCL [26] | - | 71.0 | |
| Weakly | ICCV 21’ | RTFM [13] | 77.8 | 84.3 |
| AAAI 22’ | MSL [43] | 78.6 | 85.6 | |
| ECCV 22’ | S3R [98] | 80.3 | 85.9 | |
| ECCV 22’ | CSL-TAL [97] | 71.7 | - | |
| CVPR 22’ | BN-SVP [105] | - | 83.4 | |
| TCSVT 23’ | Yang et al. [100] | 77.7 | 81.5 | |
| AAAI 23’ | UR-DMU [103] | 81.7 | 87.0 | |
| CVPR 23’ | ECUPL [102] | 81.4 | 86.2 | |
| CVPR 23’ | CMRL [107] | 81.3 | 86.1 | |
| CVPR 23’ | TEVAD [104] | 79.8 | 84.9 | |
| AAAI 23’ | MGFN [99] | 80.1 | - | |
| CVPR 23’ | UMIL [101] | - | 86.7 | |
| ours | LAP | 82.6 | 87.7 |
@@
In this section, the performance of our LAP model is evaluated on four datasets, namely XD-Violence [87], UCF-Crime [96], TAD [106] and ShanghaiTech [95]. The area under the precision-recall curve, also known as the average precision (AP) is employed as the evaluation metric for XD-Violence following the protocol in [102]. For UCF-Crime, TAD and ShanghaiTech, the area under the curve (AUC) of the frame-level receiver operating characteristics (ROC) is used instead. Specifically, AUC represents the AUC for all videos, while AUC focuses only on abnormal videos. The false alarm rates for all videos (FAR) and abnormal videos (FAR) are also reported in our ablation studies.
| Type | Source | Method | TAD | ST |
| AUC(%) | AUC(%) | |||
| Semi | ICCV 17’ | Luo et al. [64] | 57.9 | - |
| CVPR 18’ | Liu et al. [65] | 69.1 | 72.8 | |
| Weakly | CVPR 21’ | MIST [85] | 89.2 | 94.8 |
| ICCV 21’ | RTFM [13] | 89.6 | 97.2 | |
| TIP 21’ | WSAL [106] | 89.6 | - | |
| CVPR 23’ | ECUPL [102] | 91.6 | - | |
| CVPR 23’ | UMIL [101] | 92.9 | 96.8 | |
| CVPR 23’ | CMRL [107] | - | 97.6 | |
| ours | LAP | 93.1 | 97.4 |
XD-Violence [87] is a multi-scene public dataset for VAD. It consists of a total duration of 217 hours and includes 4,754 untrimmed videos. The training set contains 3,954 videos while the test set comprises 800 videos. XD-Violence covers various unusual types of events including abuse incidents, car accidents, explosions, fights, riots, and shootings.
@@
UCF-Crime dataset [96] is a large-scale collection of 1,900 videos captured by surveillance cameras in various indoor and outdoor scenarios. This dataset consists of 1,610 labeled training videos and 290 labeled test videos with a total duration of 217 hours. The dataset covers 13 types of anomalous events such as abuse, robbery, shootings and arson.
TAD [106] is a dataset for anomaly detection in traffic scenes, consisting of 400 training videos and 100 test videos, with a total of 25 hours of video footage. It covers seven types of real-world anomalies.
ShanghaiTech consists of surveillance videos from different scenes on a campus [70]. The training set contains 237 videos while the testing set has 200 videos.
| Baseline | FS | MPL | PAL | UCF | XD | TAD | ||

### v2

# Learning Suspected Anomalies from Event Prompts for Video Anomaly Detection

@@ Abstract
Most models for weakly supervised video anomaly detection (WS-VAD) rely on multiple instance learning, aiming to distinguish normal and abnormal snippets without specifying the type of anomaly. However, the ambiguous nature of anomaly definitions across contexts may introduce inaccuracy in discriminating abnormal and normal events. To show the model what is anomalous, a novel framework is proposed to guide the learning of suspected anomalies from event prompts. Given a textual prompt dictionary of potential anomaly events and the captions generated from anomaly videos, the semantic anomaly similarity between them could be calculated to identify the suspected events for each video snippet. It enables a new multi-prompt learning process to constrain the visual-semantic features across all videos, as well as provides a new way to label pseudo anomalies for self-training. To demonstrate its effectiveness, comprehensive experiments and detailed ablation studies are conducted on four datasets, namely XD-Violence, UCF-Crime, TAD, and ShanghaiTech. Our proposed model outperforms most state-of-the-art methods in terms of AP or AUC (86.5%, 90.4%, 94.4%, and 97.4%). Furthermore, it shows promising performance in open-set and cross-dataset cases. The data, code, and models can be found at: https://github.com/shiwoaz/lap.
@@
Weakly Supervised, Video Anomaly Detection, Event Prompt, Multiple Instanse Learning.I Introduction
Video anomaly detection (VAD) [1, 2, 3] is crucial in video surveillance, given the extensive use of surveillance cameras. The task of VAD is to determine whether each frame in a video is normal or abnormal, which poses a significant challenge as it is not feasible to train a model with complete supervision. Consequently, weakly supervised learning methods (WS-VAD) [4, 5, 6] that solely rely on video-level annotations have gained importance and popularity in recent years.
The general paradigm of these methods involves utilizing convolutional networks such as 3D ConvNet (C3D) [7], inflated 3D ConvNet (I3D) [8], or vision transformer [9] to extract visual features and aggregate spatio-temporal information between consecutive frames. Subsequently, an anomaly-detection network is trained using multiple instance learning (MIL) [10]. This approach simultaneously maximizes and minimizes the top- highest scores from individual anomaly and normal videos, respectively. Most methods [10, 11] only focused on the visual-related modality, while some [12, 13] have incorporated semantic descriptions into videos. However, such semantic information was simply fused with the visual one, instead of delving into the underlying meaning of the textual descriptions. As a result, the MIL based approaches often suffer from a relatively high false alarm rate (FAR) and low accuracy in detecting ambiguous abnormal events.
@@
The main contributions of this work are threefold:
The new textual prompts describing the abnormal events are introduced into weakly supervised video anomaly detection. Giving the explanation of what is anomalous, the score predictor can implicitly learn more details about the anomalies. It leads to incredible performance on open-set and cross-database problems.
A new multi-prompt learning strategy is proposed to provide an overall understanding of normal and abnormal patterns across different videos, while MIL is limited to individual videos.
@@ II-A Weakly Supervised Video Anomaly Detection
The weakly supervised methods tackle frame-level anomaly detection by video-level annotations. Most of them are based primarily on multiple instance learning (MIL) due to limited annotated labels [10]. However, conventional MIL faces challenges in providing sufficient supervision for various anomalies, leading to misclassifications and a high false alarm rate. To address these issues, Yu et al. propose cross-epoch learning (XEL) [4], which stores hard instances from previous epochs to optimize the anomaly predictor in the latest epoch. Additionally, dual memory units with uncertainty regulation (UR-DMU) [16] extend the anomaly memory unit into learnable dual memory units to alleviate the high false alarm rate issue. Another approach called robust temporal feature magnitude learning (RTFM) [11] trains a feature magnitude learning function to effectively recognize positive instances. All these methods are based on single or multiple visual modalities, including RGB and optical flow.
@@ II-B Prompt Tuning for Visual Tasks
In the realm of pre-trained foundation multimodality models, a cost-effective prompt tuning approach is gaining traction for adapting models to downstream tasks in the domains of natural language processing (NLP) [17, 18] and computer vision [19].
The concept of prompt tuning originated in computer vision to tackle zero-shot or few-shot image tasks by incorporating semantic guidance. Multimodal models such as CLIP [15] leverage textual prompts for image classification, demonstrating state-of-the-art performance. In the video domain, Sato [20] explores prompt tuning for zero-shot anomaly action recognition, using skeleton features and text embeddings in a shared space to refine decision boundaries. Wang et al. introduce prompt learning for action recognition (PLAR) [21], which incorporates optical flow and learnable prompts to acquire input-invariant knowledge from a prompt expert dictionary and input-specific knowledge based on the data.
A previous effort, the prompt-based feature mapping framework (PFMF) [22], applies prompt-based learning to semi-supervised video anomaly detection. PFMF generates anomaly prompts by concatenating anomaly vectors from virtual datasets and scene vectors from real datasets, guiding the feature mapping network. However, the prompt in PFMF defines anomalies at the visual level, introducing ambiguity. In our work, we propose textual anomaly prompts based on prior knowledge to mine fine-grained anomalies to achieve high performance.
Several recent studies have introduced additional information or tasks to maximize the CLIP’s effectiveness in WS-VAD. CLIP-assisted temporal self-attention (CLIP-TSA) [23] incorporates temporal information into CLIP features using a self-attention mechanism. In contrast, VadCLIP [13] delves deeper into aligning textual category labels with CLIP’s visual features to enhance its WS-VAD performance. Unlike VadCLIP, which constructs learnable prompts based on class labels, our approach designs event prompts to describe specific anomaly-related situations, eliminating the need for additional supervised information or learning tasks.
III Methodology
@@ III-A Feature Synthesis
To clarify, we acquire the visual features and from video snippets in the abnormal and normal bags, utilizing the visual encoder of a CLIP model [15]. Given that many VAD videos, primarily from surveillance, often lack associated text descriptions, we leverage a pre-trained visual-to-text encoder from SwinBERT [25], following TEVAD [12], to generate descriptions for each video snippet. These textual descriptions then undergo processing by the semantic feature extractor (SimCSE [26]), producing corresponding semantic features and for abnormal and normal video snippets. With extracted visual features and semantic features in hand, we feed these features into a multi-scale temporal network (MTN) to obtain both local and global temporal fused features.
@@
where symbolizes a feature alignment and fusion operation. It can be either a concatenation or addition. Subsequently, the anomaly scores and can be calculated by applying and to a score predictor. Typically, this predictor takes the form of a multi-layer perceptron (MLP) [27], expressed as:
@@ III-B Multi-Prompt Learning
From Equation 5, it can be seen that the top- strategy focuses only on a few snippets with the highest scores within an individual video. Moreover, the top anomaly scores in an abnormal video may not be from an abnormal snippet. Therefore, a new textual prompt dictionary consisting of anomaly prompts is designed to link abnormal video snippets from different videos. Unlike the category annotations used in VadCLIP [13], we expanded the single word annotations into complete anomaly sentences, like “someone is doing something to whom” or “something is what”. These sentences can better describe the events/actions related to a certain anomaly category. Then, the prompt dictionary is constructed as a set of these anomaly sentences, e.g. “A man is shooting a gun” or “Something is on fire”. As depicted in Figure 2, these prompts undergo the same semantic extraction process (SimCSE [26]) as the earlier video captions, generating their respective semantic features . Subsequently, we calculate the similarity between each prompt in the dictionary and every snippet in to construct an anomaly matrix as,
in which denotes the -norm. The consideration of is dismissed here since there are no abnormal snippets in the normal bag. In essence, each element in provides insights into the probable type of anomaly associated with each snippet or indicates where a predetermined abnormal event might occur. Figure 3 offers a visual representation of , where frames containing abnormal events exhibit more pronounced colors. Notably, there is a discernible alignment between the frames and prompts.
@@ III-C Pseudo Anomaly Labeling
where and are the -th element of p and c, mean and std are the mean and standard deviation considering the anomaly vector c, and is a hyper-parameter. Then, the anomaly score predictor can be trained in a fully supervised manner, through a pseudo anomaly loss ,
@@
where and are hyper-parameters utilized in our model. Importantly, it’s worth noting that the MPL and PAL modules are trained collaboratively. During the inference stage, the test samples will only traverse the feature extractors and the predictor to acquire abnormal scores, and the MPL and PAL modules incur no additional computational cost.
III-D Inference Process
The inference process is identical to the baseline model, i.e. TEVAD [12], which is the left part of Figure 2 without the prompt dictionary. We initially extract visual and text features, which are processed through the feature alignment and fusion operation. Then, the fused features are fed into the anomaly predictor to calculate the anomaly score for each video snippet.
@@ IV Experiments
| Type | Source | Method | Feat. | XD | UCF | |
| AP | AUC | AUC | ||||
| all | abn | |||||
| Semi | CVPR 16’ | Conv-AE [29] | AE | 30.7 | - | - |
| CVPR 22’ | GCL [30] | CNN | - | 71.0 | - | |
| Weakly | ICCV 21’ | RTFM [11] | CLIP | 78.3 | 85.7 | 63.9 |
| AAAI 22’ | MSL [31] | ViT | 78.6 | 85.6 | - | |
| ECCV 22’ | CSL-TAL [32] | I3D | 71.7 | - | - | |
| CVPR 22’ | BN-SVP [33] | I3D | - | 83.4 | - | |
| CSVT 23’ | Yang [34] | I3D | 77.7 | 81.5 | - | |
| AAAI 23’ | UR-DMU [16] | CLIP | 82.4 | 86.7 | 68.6 | |
| CVPR 23’ | ECUPL [35] | I3D | 81.4 | 86.2 | - | |
| CVPR 23’ | CMRL [36] | I3D | 81.3 | 86.1 | - | |
| CVPR 23’ | TEVAD [12] | I3D | 79.8 | 84.9 | - | |
| AAAI 23’ | MGFN [37] | ViT | 80.1 | - | - | |
| CVPR 23’ | UMIL [28] | XCLIP | - | 86.7 | 68.7 | |
| ICIP 23’ | CLIP-TSA [23] | CLIP | 82.2 | 87.6 | - | |
| CVPR 24’ | Wu et al. [38] | CLIP | 66.5 | 86.4 | - | |
| AAAI 24’ | VadCLIP [13] | CLIP | 84.5 | 88.0 | 70.2 | |
| ours | LAP | CLIP | 86.5 | 88.9 | 73.0 | |
| ours | LAP | CLIP | - | 90.4 | 76.1 | |
@@
In this section, the performance of our LAP model is evaluated on four datasets, namely XD-Violence [39], UCF-Crime [10], TAD [40] and ShanghaiTech [41]. The area under the precision-recall curve, also known as the average precision (AP) is employed as the evaluation metric for XD-Violence following the protocol in [35]. For UCF-Crime, TAD and ShanghaiTech, the area under the curve (AUC) of the frame-level receiver operating characteristics (ROC) is used instead. Specifically, AUCall represents the AUC for all testing videos, while AUCabn focuses only on abnormal videos in test set. The false alarm rates for all videos (FARall) and abnormal videos (FARabn) are also reported in our ablation studies.
IV-A Datasets
@@
XD-Violence [39] is a multi-scene public dataset for VAD. It consists of a total duration of 217 hours and includes 4,754 untrimmed videos. The training set contains 3,954 videos while the test set comprises 800 videos. XD-Violence covers various unusual types of events including abuse incidents, car accidents, explosions, fights, riots, and shootings. UCF-Crime dataset [10] is a large-scale collection of 1,900 videos captured by surveillance cameras in various indoor and outdoor scenarios. This dataset consists of 1,610 labeled training videos and 290 labeled test videos with a total duration of 217 hours. The dataset covers 13 types of anomalous events such as abuse, robbery, shootings and arson. TAD [40] is a dataset for anomaly detection in traffic scenes, consisting of 400 training videos and 100 test videos, with a total of 25 hours of video footage. It covers seven types of real-world anomalies. ShanghaiTech consists of surveillance videos from different scenes on a campus [42]. The training set contains 237 videos while the testing set has 200 videos.
IV-B Implementation Details
The dimension of the visual features extracted by CLIP(ViT-L/14) [15] is 768, while the dimension of semantic features is also 768. The prompt dictionary capacity is set to 30 for UCF-Crime, XD-Violence and TAD datasets, 25 for ShanghaiTech. The batch size is set to 64 for TAD dataset, and it is halved to 32 on the other three datasets. The number of snippets per video is set to 64 for all datasets. The feature operation is set as, a) concatenation for UCF-Crime, b) addition for the other three datasets. The hyper-parameters , , and are consistent across all datasets. The Adam optimizer is utilized with a learning rate of 0.001 and weight decay of 0.005 during the training process.
| Type | Source | Method | Feat. | TAD | ST |
| AUC | AUC | ||||
| Semi | ICCV 17’ | Luo et al. [43] | - | 57.9 | - |
| CVPR 18’ | Liu et al. [44] | - | 69.1 | 72.8 | |
| Weakly | CVPR 21’ | MIST [45] | UNet | 89.2 | 94.8 |
| ICCV 21’ | RTFM [11] | I3D | 89.6 | 97.2 | |
| TIP 21’ | WSAL [40] | I3D | 89.6 | - | |
| CVPR 23’ | ECUPL [35] | I3D | 91.6 | - | |
| CVPR 23’ | CMRL [36] | I3D | - | 97.6 | |
| CVPR 23’ | TEVAD [12] | CLIP | 92.3 | 97.3 | |
| CVPR 23’ | UMIL [28] | XCLIP | 92.9 | 96.8 | |
| ours | LAP | CLIP | 94.4 | 97.4 |
IV-C Comparison Results
Quantitative analysis. The comparisons between our LAP model and other state-of-the-art (SOTA) WS-VAD models on the XD-Violence and UCF-Crime datasets are presented in Table I, and Table II shows the comparisons on the TAD and Shanghaitech datasets. It can be seen that the proposed model outperforms almost all the other methods in all datasets.
Specifically, our model achieves the highest AP of 86.5% on the XD-Violence dataset, outperforming the second best method VadCLIP [13] by 2.0%, which also combines RGB and text data. Unlike the single-word description, i.e. class labels, used in VadCLIP, the event-level descriptions in our LAP model can provide much richer information, leading to a better understanding of the anomalies. Another ClIP-based method (CLIP-TSA [23]) leverages the visual features from CLIP (VIT/B), while a transformer is employed to enhance its features. However, due to the lack of semantic guidance, its AP falls 4.3% below the proposed LAP. The performance of the other compared the methods are also limited by the absence of efficient anomaly definitions.
In the UCF-Crime dataset, our LAP model achieves an AUCall of 88.9%, surpassing the most recent methods by at least 0.9%, including VadCLIP [13] (88.0%), CLIP-TSA [23] (87.6%) and UMIL [12] (86.7%). Notably, the learnable prompts of VadCLIP are based on class labels, which leads to a video-level anomaly matching. While our concise descriptions of basic suspected anomaly events can effectively match the segment-level features. This difference results in a relatively high AUCabn of our LAP (73.0%) comparing to VadCLIP (70.2%). It is important to note that if we use more accurate text descriptions [46] of each snippet, we can achieve a higher AUCall of 90.4% and an AUCabn of 76.1%. The results indicate that our model can effectively utilize the textual prompts of abnormal events for an accurate detection.
Table II shows the comparisons on the other two less challenging datasets. The AUC of our approach (94.4%) is constantly higher than all SOTA methods compared [35, 40, 28] by a margin of 1.5% to 4.8% on TAD dataset. And our method achieves the second highest AUC (97.4%) in ShanghaiTech, which is only 0.2% lower than the best CMRL method [36]. Noting that, if we switch our visual extractor from CLIP to I3D [8] as the same as CMRL, the AUC will be boosted to 98.0% (0.4% higher CMRL). It indicates that the ShanghaiTech dataset is relatively less complex, while I3D [8] is good enough. The details will be discussed in Section IV-E. For fair comparation, we re-implement TEVAD[12] with visual features from CLIP. The AUC of our LAP exceeds TEVAD for 2.1% in the TAD dataset and 0.1% in the ShanghaiTech dataset. This minor performance gain on ShanghaiTech is due to the anomalies on campus being actually common activities such as riding, skating, and driving on the road, which are quite different from our suspected anomaly descriptions such as fighting, firing, or clashing.
Overall, these results highlight the superior performance of our LAP model compared to state-of-the-art methods on all four datasets in terms of both AP and AUC metrics. For fair comparison, the UMIL [28], TEVAD [12], CLIP-TSA [23], VadCLIP [13], and the work by Wu et al. [38] are based on the same feature extractor (CLIP) as our LAP.
Qualitative analysis. To further demonstrate the effectiveness of our method, the qualitative comparisons between our approach (LAP) and the TEVAD SOTA method [12] are visualized in Figure 4. The normal and abnormal frames of videos from the UCF-Crime and XD-Violence datasets are presented along with their corresponding frame-level anomaly scores, while green and red dashed rectangles indicate normal and abnormal ones, respectively. As shown in the figures, our method not only outperforms TEVAD [12] in terms of anomaly detection ability but also reduces false alarms on normal parts.
Our prompt dictionary contains event descriptions for various conditions. For the UCF-Crime social dataset and the TAD traffic dataset, Figure 5 illustrates the distributions of matched suspected anomalies in the anomaly vector c, showcasing how our prompt dictionary operates. Since the majority of anomalies in UCF-Crime are linked to human behavior like fights, robbery and violence, the predominant subject is “person” and the most frequent activities include falling and using weapons. While in the other circumstance, the Traffic Anomaly Dataset (TAD) is consists of anomalies caused by traffic accidents. As expected, “car” is the main subject, and the most common activities involve smoking and crashes. It indicates the effectiveness of our proposed event prompts.
IV-D Ablation Studies
Components. The proposed prompt-related components, i.e. feature synthesis (FS), multi-prompt learning (MPL) and pseudo anomaly labeling (PAL) are the keys to our superior performance in VAD. The ablation results of these three components on three datasets are shown in Table III. The baseline module is a visual-only branch MIL-based network with a CLIP feature extractor [15]. By cooperating text branch for feature synthesis, our model achieves 1.2%, 1.8% and 0.4% AUCall improvement, respectively, which shows the efficiency of semantic information. The MPL module can also improve the AUCall for all datasets by 0.3%, 0.4% and 0.1%, while the AP on XD-Violence and AUCabn on UCF-Crime are also boosted by 0.9% and 0.3%. Further incorporating the PAL module yields better results. It outperforms the baseline by 1.9% in AUCall and 6.0% in AUCabn on UCF-Crime, 0.7% in AUCall on TAD, as well as 2.4% in AUCall and 5.2% in AP on XD-Violence.
Prompt format. The prompt format plays an important role in the proposed LAP model. Thus, two different formats are tested in this experiment. One is organized by anomaly phrases, such as ”falling down” or ”on fire”. The other contains complete anomaly sentences, like ”someone is doing something to whom” or ”something is what”. As shown in Figure 6a, the sentence-based prompt dictionary outperforms the phrase-based one by 3.1% on XD-Violence and 0.7% on UCF-Crime, respectively. It suggests that prompts containing richer information are more helpful in identifying suspected anomalies.
Pseudo anomaly threshold. As required by the PAL module, pseudo labels are determined according to the threshold . The dynamic threshold given in Eq. 13 is used in previous experiments, which is based on the distribution of the data in the current batch. The hyper-parameter will determine the number for pseudo anomalies. When it is set to 0.5, 1.0 and 2.0, the AUC results on UCF-Crime are 88.05%, 88.90%, and 88.21%, correspondingly. Another static threshold strategy is also compared in this test, while is set to 0.5 as prior knowledge. As shown in Figure 6b, the dynamic threshold is better than the static one, whose AP and AUC are 2.1% and 0.5% higher on XD-Violence and UCF-Crime datasets.
| Baseline | FS | MPL | PAL | UCF-Crime | XD-Violence | TAD | ||